The APOcrypha

Volume 80, Issue 6

Spring 1994ExecComm

President  Dave Cho  x5-6603  dcctdw
SVP  Bridget Spitznagel x5-6152  baspitz
MVP  Barbara Manganis x5-7407  autumn
AVP  Kate Curtis  628-5384 kcurtis
Treasurer  Kevin Dahm  628-5384 kdahm
Fellowship  Ann Yen  x5-7672  aiyen
APOcEd  Amy Hsu  x5-7678  aihsu
ASC  Traci Williams  x5-8455  tlw
PubDir  LeaAnn Collins  x5-8526  troth
Historian  Dima Dribinsky  x5-7413  dimon

Chair, Advisory Comm. Joe Kuchta

Brotherhood Expectations

Attend at least 6 service projects, including 2 on-campus and 2 off-campus Attend at least 2 fellowship events, including one of the major ones Attend at least 5 Chapter Meetings and 1 ExecComm Learn, teach, or practice a chapter skill Hold a leadership position within the chapter (e.g. PC, sub-PC, hold office) Attend the Pledge Term Review Meeting Attend ceremonies

Pledge Requirements

Attend at least 6 service projects, including 2 on-campus and 2 off-campus Attend at least 2 fellowship events, including one of the major ones Attend at least 2 Chapter Meetings and 1 ExecComm Attend at least two-thirds of scheduled Pledge Meetings Learn or teach a chapter skill Get a Big Brother within two weeks of pledging PC or sub-PC a service project or fellowship event Attend the Pledge Feedback Meeting

The APOcrypha

Volume 80, Issue 6 Wednesday, April 20, 1994

Sometime between Thursday night and Friday morning, $97 disappeared from the strongbox, along with $6 from the treasurer’s drawer. There are lots of ways this could have happened; the key to the strong box was in the treasurer’s drawer, which is not secure.

Recommendations: Make sure the office is always secure and that the door is locked when there is no one conscious in the front part of the office. The missing money will be replaced from the executive discretionary fund so that Planned Parenthood (Big Screw) does not lose out. All receipts and money should now be deposited into the strongbox. No money should go into the treasurer’s drawer. The keys for the strongbox will be kept in the safe. A report was filed with the CPs. Possibilities: Talk to CAC about getting a retractor on the door. Repin the door lock so that it is more secure, but depends on how MIT feels about this.

PFTGOTOs:

Barbara: Richard and I have EMT exams on the 30th. Hannah: Premiere Orchestra concert, 2 pm, Saturday the 23rd. Dave: Drop Date Concert Choir concert, Verdi’s “Requiem,” Friday the 22nd, 8 pm in Kresge. Barbara: 6.111 still sucks. Dave: 6.003 sucks.

More Spring Breaking

Wait! Wait! Amy needs another 3 pages of material!

So I’ll write what I did during Spring Break.

Easy.

Friday, I did something, I’m sure. I think I slept.

Saturday, I think I had a netrek game.

Sunday, I think a bunch of us practiced ballroom dancing. Ballroom dancing is way cool. There are a bunch of different forms with totally different mindsets. Swing is a really casual and informal form. Rumba is a very sultry form; it’s not meant for seducing people, but them becoming a part of the chapter more quickly. They don’t have office keys so they can’t hang out in the office, unless, of course, lots of brothers are already there. :) At CPPC we discussed communication channels between the chapter and the pledges, or, rather, lack thereof. I hope to help the MVP make the improvements suggested at CPPC a reality. The idea is to make sure that the pledges feel a part of the APO community as quickly as possible and to make them feel more “clued-in.”

Chapter Skills: Holding another chapter skills weekend is an option. I’d certainly like to continue the push to revive silkscreening and train press ops to the point where the skills have enough momentum to keep themselves going. I would like to see both the silkscreen room and the press shop used extensively over the summer.

$ Money $: The first thing I’d like to do is have someone (maybe me, maybe not) go talk to the registrar’s office. Maybe we’ll get some useful information like what they see happening to Reg Day in the next few years. Then we need to analyse what we are spending money on (sounds like a job for the new treasurer) and figure out where we could best curtail our spending. We spend about $4000 a year; do you know what we spend it on?

Inter-Chapter Stuff: Yeahh for Sectional Service Project. Maybe we could organise a service project between all the chapters in Boston; it shouldn’t be too hard. Conclave can be fun; come and keep me company there.

Academics: Always come first. If anyone has an idea of a thesis topic for me, preferably one that has something to do with databases and multi-media educational initiatives/tools, please let me know.

Fellowship: PARTY!! Well, not just that, that and much more. I like study breaks. I like some of the more innovative breaks that I’ve heard of in the past, like study breaks in elevators with popcorn. Spontaneous fellowship is always good; be spontaneous. The fellowship program is also a way of making the pledges feel more at home.

Nationals: Dallas in December. Try forgetting about how much it costs to fly to Dallas and thinking instead about whether you would like to represent the chapter at the national convention. Do you want see first hand what happens to the motion (about blindfolding

2 That’s it for this term’s CPPC. Many thanks to everyone who came and participated; literally speaking, it couldn’t have happened without you.

Now for the hard part: now for implementing the changes. I wish the chapter and imminently-incoming ExecComm luck with their plans.

dave President

Stuff from mrn: Random News on AX Alums

[Editor’s Note: mrn is our APO Regional Director.]

Recently spotted in Washington, at the rechartering of Central Washington University: Liz Wise and Jock Young, AX circa mid-70s. Jock is an advisor to the new chapter.

And closer to home, Rick Hartley, an AX alum, was elected as Section 99 (New Jersey) Chair this past weekend.

Alpha Chi must be doing something right in the leadership development department. This means every Section in Region I has been chaired or represented by AX alums (and of course, there’s your beloved Regional Director, too. ;) Not to mention Sections in other parts of the country, most notably Region X, and in fact, the participation at all levels of the Fraternity.

mrn

**************

In Memoriam: Earle M. Herbert

It is my sad duty to report that past National President Earle Herbert passed away this past Friday.

To be honest, I am somewhat at a loss for the right words here. Suffice it to say that Earle will be greatly missed, and I am just so happy that my life was touched by this great man. Alpha Phi Omega be, but should not be relied upon entirely. The service program itself would have to grow in some direction (either larger projects, or more of them) to maintain current standards of per-brother service, and the recruitment and membership retention programs would have to grow to maintain the size of the larger chapter. (Actually, there is currently no membership retention program.) As for the minor issue of an increased number of people running for offices: In the (fairly recent) past, elections consisted of finding a candidate for each office and electing her/him. Now that we have a greater opportunity for choice, we will have to decide what criteria are actually appropriate (for example, should one person be favored because this is the last term she/he can reasonably be expected to hold this office? should… never mind, you can come up with weird examples on your own). An increased size may affect leadership in other ways as well (e.g., people may become more project-specialized, or officers might run projects only rarely.)

Our relations with other groups are likely to change as well. We are beginning good relations with the chapter at BU; there is another service group (Circle K, largely Baker-based) on campus which has expressed interest in establishing occasional relations; people are trying to create the tradition of a Sectional Project; more people in the chapter have started attending conclaves (yay). The chapter’s selfimage (how members perceive AX itself) is likely to affect the way that we see and interact with (or don’t interact with) outside groups.

So what am I doing next term, anyway?

(This new VI-3 degree program is confusing :-S. I used to have everything planned out… then they changed everything. Cool! techinfo has the course catalog online! … but not degree requirements.) My academic commitments will be fairly serious (though nothing to really worry about). Call it Japanese III for a HASS-D, and two or three CS classes (depending on which ones they are.)

Non-academic commitments: If I am elected to a major office, I will not participate in any organized musical endeavors. (They seem to be specially planned to conflict with APO.) It is unlikely that I will be able to find a paying UROP. If given the opportunity (the fall class is smaller, so they don’t hire as many) I will lab-assist 6.001 again; that takes about eight hours a week arranged any way I want. Thus, I should be able to attend all meetings (and open meetings,

6

CPPC Report

I want to start with something that I feel is really important.

After CPPC, quite a number of people came up to me and said something like “Great meeting” or “That was an excellent CPPC” or compliments along those lines.

Not wanting to sound either unappreciative or overly modest, but I really must respond via:

“Bulls—.”

Let’s get real.

The President does not make the CPPC. Just because I was moderating the meeting does not mean the CPPC was “good” or “better” or anything. Now perhaps I moderated it well. But this wasn’t anyone’s point.

I feel, and I think others would agree, that this CPPC was excellent in identifying problems and starting on solutions. This is good. But this is not because of me.

It’s because of the people who were there, who put in three hours to stay awake, to concentrate on what people were saying, to listen, to think, to try.

CPPC’s success emanated from two things: that a lot of people were there and that those people tried. This is the case with most things in the chapter. Never forget this. Never stop trying. Never stop trying to work with new people. Never stop trying new ideas.

So. Onwards to the rest of my report.

In attendance: Elth Ogston, Hannah Jang, Cris Lin, Junko Tamiya, Charley Hamilton, Ping Huang, Yale Zussman, Kate Curtis, Alan Meisler, Jeffrey Poore, Richard Barbalace, Ezra Erb, bert Dvornik, Cathy Coury, Glenn Berry, Ann Yen, Henning Colsman-Freyberger, Kevin Dahm, dave cho.

Time: 8:45 pm to 12:05 am.

The topics brought up for discussion were • Pledge Manuals at Pledge Ceremony Receptions (see CPPC

report).

• Pledge Study Break, preferably early in the term. Get them to organize a study break. (Ooooh, ahhh, really hard there.) The point is not a leadership exercise (except for the PC), but rather, an opportunity for pledges to meet brothers and vice versa. In addition, to also get used to working together: a not-so-subtle setup for Pledge Project.

• I think it’s policy, but if it isn’t, I want the Pledge and Voting policy to be voted on at I-Banquet.

How would I accomplish my goals? The Big Brother Network doesn’t work; it’s too big. Doing it along would be a lose. Thus, I’d be getting a MemComm. It’d be á la ServComm. Three or four people. I’d be dispatching events as they came in and making final decisions, but otherwise, working closely with these people in keeping in touch with the pledges. Now, obviously, if all the Big Brothers are angelic, getting eight hours of sleep a night, their weekly netrek game, and three balanced meals daily, things will obviously be OK. But if they don’t, well, that’s why I want to be MVP, so i can play their netrek games for them.

Where would I hold events? What a ludicrously stupid question. It was obviously constructed to have a “where” question.

Closing comments:

1. Yes, I know about the time commitments.

2. Yes, I can be serious. But I think humor is a good thing for an MVP, and while not everyone can appreciate a netrek joke, it still is a dumb joke, and as we all know, dumb jokes are always the not-quite-right form of humor for any situation.

3. I really want this job, so don’t ask me that.

4. I know I’ve missed a lot of important information, but since I won’t be at a 6.004 test tomorrow night, you can just ask me.

dave dave cho MVP candidate

10 CPPC. I know that whatever I said there made people laugh (of course there was a lot of that going on). What I basically talked about was this: I am an Athena nerd (and God help me but I know that I am not alone). This means that I read my mail almost three or four times most days. However, I listened to what was going on in the Chapter BEFORE I pledged. That was how I got a lot of my info. I listened. I didn’t have to ask much. I just knew there was an apo-news mailing list, and I added myself. (I do think I remember asking Barbara and her telling me about the list.) I read the mail, sometimes with no clue what the main point was. But sometimes, I got a lot of info out of the messages. Like what APO is about. What we do. And of course, I pledged, and then everything that came in over apo-news was informative. Now I am an Athena nerd, remember, so this method works for me. I would say for those non-Athena nerds, we need to find a way to convey this info to them. (I am not going to elaborate because this was a big subject in CPPC, and Dave will probably fill a lot of you in.) The other thing I mentioned was some thoughts I have about apo. People might think I am sweet-talking the chapter a little, but I am not being untruthful when I say that APO has grown on me… I have met a whole large number of new friends through APO. I have found a place that I like to hang out. In fact, doing homework at the office is somewhat more convenient especially since my computer is far away… PCing chapter weekend is another report that I will write, and I am sure that you will all see this later. But I must say that I had a good time PCing. Granted, we will do a tool list next time so that we don’t end up with extra tools. But it looked like everyone enjoyed the weekend and I know I did. I must now put in a good word for my fellow pledges. I must say that I still can not believe we have 18 pledges. But a lot of them I do know, and I must say that there are some really great people out there. I would like to thank my PCs for a job well-done: Cathy, Transport PC; Aphrodite, Grant, and Hannah, Food PCs; Junko, who survived her second PCship as my Manpower PC, and Cathy as temporary (Sunday) on-site PC. Thanks to all the pledges and brothers who came to help at CWII. Thanks to everyone for making me feel welcome in APO. I wish everyone luck in the final weeks and of course, good luck to APO office candidates

YiLFS & Pledging, Jeffrey myself from those organizations now that I am a graduate student. There is probably about an even chance that I will be signed up as an associate advisor.

I believe I could do a good job as APOcrypha editor, and it is a job which I believe I would enjoy and have considered running for several times before. I know I can handle the mechanics part of being APOcrypha editor: I am certainly familiar enough with Athena and LaTEX, and as a past AVP I know how to deal with labels, mailings, etc.

However, it takes more than familiarity with the mechanics to make a good APOcrypha editor. I have been very pleased to see the APOcrypha come out on a regular basis this term, and my intent if elected will be continue to put out the newsletter on a regular schedule. At CPPC it was mentioned that the APOcrypha could serve as a more effective communications tool for the chapter if it were to be published several days before each chapter meeting rather than just before, so that it may serve as a reminder to brothers that there is a chapter meeting coming up, and to provide them with a copy of the agenda. I will attempt to adhere to such a schedule. I’d also like to try to work with other officers in figuring out how the chapter can be best served by the various communications media available.

In addition, I would like to see more articles published in the APOcrypha which are not minutes summaries or project reports. To that effect, I plan to target (i.e., harass) specific officers and brothers for topical articles as well as essays on issues which arise about or in Alpha Chi. Although I will not resort to threatening to publish extremist political tracts as filler (as some of you may recall past APOcrypha editors have done) in order to coerce articles from the general brotherhood, I will solicit as best I can in general as well. It is my hope to help change how people feel about the APOcrypha, such that APOcEds will not have to plead for articles in a chapter where numerous brothers will no longer feel the need to complain about lack of effective communication.

I’d also like to see more information in the APOcrypha which is less parochial in nature, i.e., more about what other service-minded students are doing on MIT campus and at other campuses/chapters. I am open to new ideas, and may try to implement some which I’ve

14 must be AWAKE and in the MAIN office, NOT in the back room where the press and the tools are.

Thank you for your cooperation.

YiLFS, Kevin Dahm APO Treasurer

In addition, at ExecComm, it was recommended that I look into getting the lock replaced (but with the same pin heights) and perhaps a spring for the door (to ensure that the door closes the last few inches).

If you have any further questions, feel free to send me mail.

CPPC

CPPC happened at Chapter Weekend, and was very successful. My report is included outside this writing.

dave President

What I did for spring break…

[Editor’s Note: This article was actually submitted before Issue 5’s deadline, but somehow it got overlooked. My apologies to Ping.]

The weather wasn’t very spring-like. And it wasn’t much of a break. But I did sleep a lot, so that was good.

Ping

Just another article. :)

It’s been a long time since I’ve been involved in an APO project/social event, and, quite frankly, it feels rather funny. (It’s Tuesday morning, just minutes before 3:00 am, so, please, bear with me. :) )

I remember being “fairly” involved in APO in my freshman year. It was really cool to go to projects and paint and/or destroy various inanimate and/or animate objects; it was really cool (though really masochistic) to manage to put together a little Semi-Formal Formal almost alone during one of Boston’s worst blizzards. And the elevation ceremony was something that is not soon forgotten…

Hmmm… why was I inactive this term? I had this belief that I could divest myself of my activities; that way, I could spend more time on my work. Now, how many of you think that actually happened? I thought so. :) There’s this little voice inside my head that keeps telling me that I would have done no better if I didn’t become inactive and managed to make the Semi-Formal Formal (sorry Gilbert… but right now my brain is so burnt out that I can’t even remember why I couldn’t make it…). But I guess no one else here has my GPA…:-/

Well, it’s 3:15. I think it’s about time to get back to my 14.02. I’m almost a half-dozen lectures behind (can you say “6.003”?), and I have an exam this Thursday on that material.

Good night, all!

Daniel

(with only \sqrt{0.2} brain cells left for tonight)

Pledge Article: EGW

Hi! It’s Emily, who has now FINALLY gotten around to writing something for the APOcrypha. Yay!!!!!! I’m a freshman. So, yes, I’m on pass/no-record, but I am still hosed. I’m taking 1.00, yes, that’s right, 1.00, the class whose professor just won the 1994 Big Screw Contest! So, if you will excuse me, I have to go back to laboring on for the ’Tute’s biggest slavedriver this year. No one can ever be done with the work for that class; Logcher was the rightful winner. AAARRRRRRGGGHHH!!!!!

YiLFS, Emily

18 I’ll be taking the next term. I will have a lot of humanities, though, because I kind of need to catch up on my HASS requirements. I might or might not also be taking 8.07. I think I’ll decide after Junior Lab ends… Without the shadow of Junior Lab looming over my head, I’ll also be able to get more active in the chapter in general.

I can’t promise that I’ll do a better job than any of the other candidates would. If I get elected, I’ll do my best. Since you are the ones who’ll have to read the APOc, I urge you to vote for whomever you think will do the best job, even if it isn’t me. =)

YiLFS and 8.14, bert =)

APOcrypha Editor: Ping Huang

Hello. I am running for the office of APOcrypha editor. Per request (or more accurately, begging^{1}) of the current APOcrypha editor, I am writing a candidate’s article. :)

I will start off by answering the standard questions that Ezra seems compelled to ask every single candidate for any chapter office, and I will be nonplussed if I get asked these questions during the elections meeting (hint).

• The previous offices I have held in the chapter are (in chronological order): Historian, PubDir, AVP, MVP, and SVP.

• I have not finalized what I will be registered for in the fall, but it will very likely consist of two graduate-level computer science classes, a women’s studies humanities class, and something random. Next term will be my ninth term as a student, and I expect to graduate in May 1995.

• My time commitments might include working part-time off campus or doing some research in an LCS or AI Lab group, in preparation for writing a thesis sometime next year. I am currently involved in the Baker Foundation and the Associate Advising Steering Committee, but I am moving to disengage

Silly Stuff

Junko and I were being silly at Chapter weekend…

I always say “and my a– is painted green” when I don’t believe her. Well, we were painting the shutters green, so she decided to paint my butt green too… so that day, my butt was TRULY painted green.

=]

Cris

Patriots Day Weekend: The Editor’s Journal

Well, here I am again, writing another article for myself. I guess I just like to read my words in print. Anyway, this past weekend (it’s now April 19, 7:27 pm was, amazingly enough, productive, relaxing, and sleepful! How is that possible? I have no idea. Time warp, maybe.

Friday: The weather was absolument fantastique! I went jogging. In the evening I watched Groundhog Day, which was pretty funny (one Kleenex, though). Anyone want to venture a guess as to how many times Phil went through that day?

Saturday: The weather was wet. I got soaked on the way to the student center for the Technique meeting. Serves me right for being late. Anyway, for three hours I helped design and paint a poster, getting my socks very dirty and pink-spotted in the process. I also spent another three hours in the darkroom doing living group reprints of Phi Kappa Sigma. Then I went to a lobster bash and stuffed myself. Oof.

Sunday: I slept late! I did three hours of work (envelope-stuffing, counting bubble sheets, tying string around old forms) for CEG, then played three hours of volleyball, then ate dinner at Café of India in Harvard Square. Hey, I don’t know how authentic it was, but it tasted good to me (onion pita bread stuff, meat samosa stuff, lamb and mushrooms stuff drowned in brown sauce, and chicken stuff drowned in orange sauce).

Monday: Slept late again! Read Book III of Paradise Lost and ac-

22 Oops, sorry. That shameless plug just happened to slip in.

1I have more to say about this later.

Why shouldn’t I be MVP? Well, I’m old. There’s something to be said for MVPs being around after their class is elevated. However, I’m not planning on disappearing from Boston. (Yes yes yes, I realize the best-laid plans never materialize, but I’ve avoided leaving Boston for six years, and things as they are certainly point to my being around Boston.) And it is true that Being Outside will preclude a great deal of personal involvement. But I hope to be at least marginally involved, and after the first two months, I don’t expect I will have a realistically significant effect.

(No, really, I’m not tired. Yes, this is a non-sequitur.)

Other reasons not to elect me. Umm… oh yes. I might be so overly sarcastic that I’ll have two pledges elevated, four continued, and eight in a mental trauma hospital. Don’t worry about it. Or that I’ll flirt with all the pledges. Umm… hmm, well, I won’t say that I’ll never flirt with them (I’m being half-facetious, half-serious here) until I at least know them (c.f., this term’s officers and pledges).

In other words, I can’t think of any other reasons not to elect me other than other superior candidates.

What would I do? I think CPPC brought up a lot of good ideas. Umm, and not wanting to be my usual superior self, some of them were ideas I had heard of and/or came up with a long time ago. (Not at all surprisingly, I had since forgotten them.)

Down to specifics:

• BBs by Pledge Ceremonies. This is tricky.

• More formalized Pledge Meetings. I think I would try to have them just before or just after ExecComm meetings. What actually goes into them would be a summer project, but it would certainly be ruled by “what do the pledges want to know and need to know at this point” more than anything else.

• I probably wouldn’t have any other pledge-specific meetings, since realistically speaking, neither they nor I have the time for them.

• Membership: recruiting

• Membership: retention

• Expansion issues: how the chapter needs to deal with its increasing size

• Relations between officers and the rest of the chapter

• Lack of PCs

Ezra’s summary of the beginning is excellent:

We started with communication between officers and brothers/pledges, with an emphasis on how to make brothers/pledges comfortable with bringing problems and difficulties to the officers. Next it mutated into a discussion of chapter meetings vs execomms, and how to make them distinct. Then it mutated into a discussion of email and email lists. From here it degenerated into trivia, and Dave wisely called a recess.

Kate’s conclusions are insightful:

Discussion about relationship between execcomm and the rest of the chapter was not really very productive except for the bits that suggested that we look at the differences between chapter meetings and execcomms and what people want from them. Short meetings are appreciated, but this should not be at the expense of discussing chapter business or ignoring business which perhaps ought to be raised.

The few concrete ideas from the beginning half of the meeting were:

1. Announce officer hours at Chapter Meetings. Also, publish them on the inside cover of the APOcrypha.

2. At Chapter Meetings, have a one-page handout of upcoming events; basically, publish a new WUAPO for Chapter Meetings.

26 whole collection from running off the side of the road into a tree. The president also makes sure the chapter stays in contact with advisors and with the rest of the world. Lately the prez also gets the job of talking about leadership at open meetings and thinking about holding leadership workshops (if I get organized, I might do more than think about it ;-) ).

What issues do I think the next president should be concerned with?

There are several changes likely to occur in the next few years (some more abruptly than others). The next president should, as always, be prepared for probable changes and be ready to deal with the problems they are likely to cause. She/he should also consider aspects of the chapter that could be changed for the better, suggest these possibilities to the relevant officers or committees, and coordinate and assist their efforts if necessary.

You may have noticed that some people in the chapter have gotten a bit nervous about the future, particularly in reference to the budget. Strictly speaking, the budget is the treasurer’s job and is usually prepared by examining the previous few budgets and deciding whether they were reasonable or not. (Note that “reasonable” and “realistic” are not necessarily related concepts. The realistic budget is a rare beast.) The general consensus among the nervous people is that the ’Tute is going to computerize Registration Day completely, thus eliminating our major source of income; however, there is no real agreement about exactly what we should do about this when it happens, nor about whether we should start doing something about it before it happens. In other treasury matters, security has recently become an increased concern.

You may also have noticed that the chapter is getting larger and that there are more people running for offices (yay :-)). These are Good Things (tm), but a good thing can still cause problems if it is not planned for. One of the immediately apparent difficulties that an increase in membership may cause is lack of communication. When there are around twenty active brothers in the chapter (roughly half of them are officers and attend meetings regularly), communication is barely an issue; word of mouth is nearly sufficient. A larger chapter, assuming some diversity in majors and living groups (e.g., not everyone lives in New House), is likely to require more organized lines of communication. E-mail makes this less of a problem than it might has lost one of its giants and a man who, more than anyone I know, epitomized our cardinal principles, especially friendship.

Earle’s memorial service will be at 1:30 pm this coming Saturday at St. John Fisher Church, 5448 Crest Road, Rancho Palos Verde, CA 90274.

Expressions of sympathy may be sent to “The Family of Earle Herbert,” 3671 Vinton Avenue, No. 8, Los Angeles, CA 90274. This is Earle’s home address, and his family will be there for several weeks taking care of personal matters.

Both Earle and his family have asked that, in lieu of flowers, donations be sent to the Alpha Phi Omega Endowment Fund, 14901 East 42nd Street, Independence, MO 64055-7347, in the name of Earle and all future pledges. From a financial standpoint, this fund is our future, and Earle always wore his pledge pin because he felt that a true brother never stopped pledging. Thus, such a gift is most fitting and would be the best honor we could give Earle, especially because of his dedication to and respect for pledges.

Personal note: I hate to ask for money. I do it with alumni because that’s part of “the job.” But I have to tell you that even $1 in Earle’s memory to the Endowment Fund will be greatly appreciated and will honor a truly great person who gave everything he had to APO. But I especially ask for your prayers for Earle and his family, and I ask you to be grateful, as I am, for the fact that APO had an Earle Herbert.

Jerry Schroeder National President

                                                                                                 gschroed@cap.gwu.edu

**************

Chapter Meeting Summary: April 6, 1994

The meeting began at 7:33 pm in the Mezzanine Lounge on the third floor of the student center. People who were there at one time or another: Dima, Ezra, Jo-Anne, LeaAnn, Henning, Dave, Bryan, Kate, Rachel Yale, Richard, Kevin, Charley, Bridget, Gilbert, Cathy,

30

Editor’s Sigh of Relief

Whew! Mucho thanQs to everyone who helped keep this issue of the APOc from being one sheet of folded paper! In fact, it probably broke several records! I may have to write an article about it! :)

YiAPOc, Amy

Candidate Articles!

President: Kate Curtis

I am running for president, because I think that I can continue to motivate the chapter, while giving ExecComm the direction and help it needs to implement the service/membership/fellowship/whatever programs for the fall. This article is really a collection of random things I’d like to see happen (not in any particular order); please let me know what you think, whether or not I get elected. :)

Summer: I’d like to see a couple of service projects and lots of random fellowship, from dinner trips to ???? And, of course, July 4th on the Esplanade. For those of you who are not going to be here for the summer, I hope that you will miss lots of fun, and we look forward to seeing you again in the fall.

Lots more service: I like going on service projects. (That’s probably why I’m still hanging around APO.) Hopefully you all do too. I’d like to ensure (or help the SVP ensure) that as our chapter grows, the number and type of projects change accordingly. We may have the manpower to run double projects some weekends (Saturday and Sunday), especially if some projects can only take up to ten or so people. I also want to ensure that the service program has a large number of service projects in the first third of the term when people have lots of free time. This will also give pledges lots of opportunities to meet brothers. To do this, the SVP is going to need an active ServComm. We need a ServComm to nurture future SVPs too. Really, it’s not that hard a job, and it’s fun. :)

Lots of pledges? Well, I hope so. Most importantly I’d like to see just for being very expressive and very passionate. Waltz, of course, is very refined and very sophisticated. In its own way, it can be seductive, but it’s the kind of head-over-heels type of feeling. Tango requires a lot of self-confidence. It’s very different from Rumba, but I can’t really put my finger on what. That leaves Samba, which is another Latin dance, and is kinda a cross between Tango and Rumba: more passion than Tango, but faster than Rumba.

Monday I tooled 6.003. Ick!

Tuesday I tooled 6.003. Ick!

Wednesday I punted by stealing as much of ExecComm as possible for dinner. Afterwards, we went and saw Schindler’s List. That is such an amazing movie… everyone should definitely see it, but do not see it alone, and don’t forget to bring Kleenex. And don’t plan on doing anything afterwards.

Thursday was the Aquarium trip. That was fun. :)

Friday, Saturday, and Sunday I tooled. And slept.

That’s all. :)

dave

More Dave

Well, that last article certainly couldn’t have pushed everything to three pages, so here’s a few random annoucements:

MIT’s Concert Choir, Chamber Chorus, and Symphony Orchestra are performing Verdi’s “Requiem” this Friday at 8 pm in Kresge Auditorium. Tickets are $2 in advance for MIT students; $10 general admission. (Sorry; the music office is being fascist.) As of Tuesday’s rehearsal, we sound fantastic. Come listen to me, Richard Barbalace, Harry Tsai, and a couple hundred other people perform. It’s really cool. Besides, when else can you see Harry Tsai in a tuxedo?

Central Connecticut State University’s (CCSU) Epsilon Delta chapter is having their semi-normal [yes, that’s right, semi-normal] on Friday, May 13th. So, if you’re like me and are going to be done with finals by then and are interested in going, talk to me. No guar-

34